OpenStudy (savy):

Find all rational roots of the equation. Give me a minute to type out what I already have.

OpenStudy (savy):

\[4x^4-x^3-2x^2-4x-72\] There will be 4 total roots; Using Descartes rule of signs; there can be 1 positive root, 3 or 1 negative root; and 0 or 2 non-real.

OpenStudy (savy):

I used p/q to figure out the possible roots (which is many so I don't want to type them all out); Used synthetic division to find out that -2 and 9/4 are roots which satisfies the 1 pos 1 neg; so how would i find out if it had non-real and what the non-real roots are?

OpenStudy (welshfella):

2 factors are (x + 2) and (9x - 4) So divide (x + 2)(9x - 4) into the equation to get another quadratic to solve.

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @welshfella 2 factors are (x + 2) and (9x - 4) So divide (x + 2)(9x - 4) into the equation to get another quadratic to solve. \(\color{#0cbb34}{\text{End of Quote}}\) If 9/4 is a root then shouldn't the factor be (4x-9) x=9/4 4x=9 4x-9=0

OpenStudy (faiqraees):

(x+2)(4x-9)(ax²+bx+c)=provided polynomial Compare coefficients and you will get a quadratic equation Use \[\large\rm x=\frac{-b±\sqrt{b^2-4ac}}{2a} \] on the quadratic equation to find out the remaining roots

OpenStudy (welshfella):

yea the factor should be (4x - 9) My mistake.

OpenStudy (savy):

-2,-2,\frac{9}{4},2i,-2i I plugged this into my answer for my homework and it says incorrect. (It says to include repeated answers thats why -2 is listed twice) I also looked in mathway to be sure I did everything correct and it said this was the answer? @faiqraees @welshfella

OpenStudy (savy):

@ganeshie8 @pooja195 @Hero

OpenStudy (welshfella):

The 2 roots -2 and 9/4 are correct. If we divide (x +2)(4x-9) into the function we get the quotient x^2 + 4 Solving x^2 + 4 = 0 gives 2 non-real roots :- 2i and -2i

OpenStudy (welshfella):

This is a quartic equation and has 4 roots not 5. -2 is not repeated.

OpenStudy (faiqraees):

Repeated roots are those roots which have a power on them when in factorized form. (Or you can say that the derivatives also has the same factor as its root)

OpenStudy (welshfella):

yes if -2 were repeated there would be a factor (x + 2)^2

OpenStudy (savy):

Ok, but it said the answer was incorrect and I have no more attempts. So, I'm confused as to what the answer was. I have my final tomorrow...